OpenStudy (anonymous):

simplify [(sec^2)t-1]/(tan^2)t

OpenStudy (anonymous):

1+tan^2(t)=sec^2(t)=> sec^2(t)-1=tan^2(t) tan^2(t)/tan^2(t)=1